OpenStudy (anonymous):

Given the acceleration, initial velocity, and initial position of a body moving along a coordinate line at time (t), a=12, v(0)=7, s(0)=-1

OpenStudy (anonymous):

what is it you want to solve for or find?

OpenStudy (anonymous):

find the body's position at time t.

OpenStudy (anonymous):

you want an equation that will tell you the position of where the body is by inputting time only, so use the kinematic equation: s = s + vt + (1/2)at^2

OpenStudy (anonymous):

they give you a=12 v(0)=7 s(0)=-1

OpenStudy (anonymous):

right, so replace the terms in the equation with what was given to you

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

oh so do i get 6t^2+7t-1

OpenStudy (anonymous):

so you have this equation: s = s + vt + (1/2)at^2 a=12 v(0)=7 s(0)=-1 so it becomes, s = -1 + 7t + (1/2)12t^2 which is s = -1 + 7t + 6t^2 or s = 6t^2 + 7t - 1 good work! ^_^

OpenStudy (anonymous):

so i gotta question so if i have the same equation but my value of a=20 cos 3t v(0)=9 s(0)=5 so how do i work that out

OpenStudy (anonymous):

How do i set it up

OpenStudy (anonymous):

just plug it in, its the same thing

OpenStudy (anonymous):

s = s + vt + (1/2)at^2 a=20 cos 3t v(0)=9 s(0)=5

OpenStudy (anonymous):

whats half of A=20 cos 3t

OpenStudy (anonymous):

oh, the 1/2 only affects the 20, so its the same as saying (1/2) * 20 the 1/2 wont affect the cos3t

OpenStudy (anonymous):

so is it 10 cos 3t^3 +9t +5

OpenStudy (anonymous):

did you get something else

OpenStudy (anonymous):

(10)cos(3t)t^2 +9t +5, it should look like this, do you see how the 3t goes with the cos and the t^2 comes from the equation I gave you